10. Use the normal distribution table, Table B.8 in Appendix B, to find the tail area associated with the following z values: b. z=-1.53 c 2-.64 d. z=4.73 t ntn Gna the teil area associated with
Answer #1

a) P(Z > 2.07) = 1 - P(Z < 2.07) = 1 - 0.9808 = 0.0192

b) P(Z < -1.53) = 0.0630

c) P(Z < -0.64) = 0.2611

d) P(Z > 4.73) = 1 - P(Z < 4.73) = 1 - 1 = 0

e) P(Z > 0.5) = 1 - P(Z < 0.5) = 1 - 0.6915 = 0.3085